OpenStudy (anonymous):

Find f' of f(x)=sin(2x)^2 at f'(pi/6)

sam (.sam.):

f(x)=sin(2x)^(2) The derivative of f(x) is equal to f'(x)=(d)/(dx) sin(2x)^(2). f'(x)=(d)/(dx) sin(2x)^(2) Remove the parentheses around the expression cos(v). (d)/(dv) sin(v)=cos(v) Simplify the derivative. (d)/(dx) sin^(2)(2x)=4sin((2x))cos((2x)) The derivative of the function is 4sin((2x))cos((2x)). f'(x)=4sin((2x))cos((2x))

sam (.sam.):

f'(pi/6)=4sin((2[pi/6]))cos((2[pi/6]))

OpenStudy (anonymous):

COuld you look at this Prove that 76^(130) + 1 is not prime

sam (.sam.):

I got to go now , you can tell others to solve that

OpenStudy (anonymous):

ok thanks though